OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x - y - 3z if x = -2, y = 1, and z = -2

OpenStudy (anonymous):

You just simply plug those numbers to the expression. 3(-2) - (1) - 3(-2) = -6 - 1 + 6 = -1

OpenStudy (anonymous):

thanks can you help me with this one ? Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

If you arrange the expression, it will look like this: 2x-2x-2y-y+5z+3z =-3y+8z